# N. Deng & Y. Wang & W. Cao

### v1

@@ 4.1 Uniform moment boundedness of the numerical solution
$\left\{\begin{aligned} &\mathrm{d}V^{\tau,N}_{t}+A_{N}^{2}V^{\tau,N}_{t}\mathrm{d}t=A_{N}e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor})\mathrm{d}t,\quad t>0,\\ &V^{\tau,N}_{0}=u_{0}^{N}.\end{aligned}\right$. (63)
Taking the inner product on (63) with $V^{\tau,N}_{t}$, we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\big\|V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}\right)=$ $-\big\|A_{N}V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}+\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor}),A_{N}\big(e^{-A_{N}^{2}(t-\lfloor t\rfloor)}-I\big)V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (64)
@@
Using Hölder’s inequality, the estimate (53) with $\gamma_{1}=\frac{1}{2}$, $\gamma_{2}=1$, and $\gamma_{3}=\frac{1}{4}$, (50) and Young’s inequality, we obtain
$\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor}),A_{N}\big(I-e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\big)V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (65)
@@
In order to estimate the third term on the right-hand side of (64), we need to establish the continuity of $U^{\tau,N}_{t}$. We rewrite (54) as
@@
Combining this with (52), we obtain
$\big\|U_{t}^{\tau,N}\big\|_{w^{1,2}_{N}}\leq$ $\left\|e^{-A_{N}^{2}(t-\lfloor t\rfloor)}U_{\lfloor t\rfloor}^{\tau,N}\right\|_{w^{1,2}_{N}}+(t-\lfloor t\rfloor)\left\|A_{N}e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\tilde{F}_{N}\big(U_{\lfloor t\rfloor}^{\tau,N}\big)\right\|_{w^{1,2}_{N}}$ (66)
@@
where $\mathscr{R}_{t}:=\big\|\mathscr{O}_{t}^{\tau,N}\big\|_{w^{1,2}_{N}}+\big\|\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}$. Utilizing (54), we can deduce
$\left\|U_{t}^{\tau,N}-U_{\lfloor t\rfloor}^{\tau,N}\right\|_{l_{N}^{2}}$ (67)
@@
By using (53) with $\gamma_{1}=0,\ \gamma_{2}=1$ and $\gamma_{3}=\frac{1}{2}$, and Assumption 2, we have
$\mathscr{K}_{1}=\left\|e^{-A_{N}^{2}\lfloor t\rfloor}\left(e^{-A_{N}^{2}(t-\lfloor t\rfloor)}-I\right)u_{0}^{N}\right\|_{l^{2}_{N}}\leq C(t-\lfloor t\rfloor)^{\frac{1}{2}}\left\|A_{N}u_{0}^{N}\right\|_{l^{2}_{N}}\leq C\tau^{\frac{1}{2}}.$ (68)
It follows from (53) with $\gamma_{1}=1,\ \gamma_{2}=\frac{1}{2}$ and $\gamma_{3}=\frac{1}{2}$ that
$\mathscr{K}_{2}\leq$ $C\int_{0}^{\lfloor t\rfloor}(\lfloor t\rfloor-\lfloor s\rfloor)^{-\frac{3}{4}}(t-\lfloor t\rfloor)^{\frac{1}{2}}e^{-\frac{1}{2}(\lfloor t\rfloor-\lfloor s\rfloor)}\left\|(-A_{N})^{\frac{1}{2}}\tilde{F}_{N}\big(U_{\lfloor s\rfloor}^{\tau,N}\big)\right\|_{l^{2}_{N}}\mathrm{d}s$ (69)
@@
Similarly, by (53) and (50) we can obtain
$\mathscr{K}_{3}\leq\frac{(t-\lfloor t\rfloor)^{\frac{3}{4}}\left(1+\big\|U_{\lfloor s\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}^{3}\right)}{1+\tau\big\|U_{\lfloor s\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}^{12}}\leq C\tau^{\frac{1}{2}}.$ (70)
By substituting (68)-(70) into (67), we obtain
$\left\|U_{s}^{M,N}-U_{\lfloor s\rfloor}^{M,N}\right\|_{l^{2}_{N}}\leq$ $C\tau^{\frac{1}{4}}+\left\|\mathscr{O}_{s}^{\tau,N}-\mathscr{O}_{\lfloor s\rfloor}^{\tau,N}\right\|_{l^{2}_{N}},$ (71)
which combined with Young’s inequality, (7), (49) and (66) yields
$\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor})-\tilde{F}_{N}(U^{\tau,N}_{t}),A_{N}V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (72)
@@
Then it follows from (64), (65) and (72)
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\big\|V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}\right)=$ $-\frac{3}{4}\big\|A_{N}V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}+C\left(1+\mathscr{R}_{t}^{8}+\tau^{-\frac{1}{3}}\left\|\mathscr{O}_{t}^{\tau,N}-\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{2}+\left\|\mathscr{O}_{t}^{\tau,N}-\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{4}\right)$ (73)
@@
Similar to the treatment of (20), we can obtain
@@
$\leq$ $C\left(1+\|u^{N}_{0}\|_{l^{\infty}_{N}}^{20}+\left\|(-A_{N})^{\frac{1}{2}}u^{N}_{0}\right\|_{l^{2}_{N}}^{10}+\mathbb{I}_{1}\big(t,3/4,\lambda_{N,1}^{2}/2,|K_{4}|^{3}+\left\|\mathscr{O}^{\tau,N}\right\|_{l^{6}_{N}}^{3}\big)+\mathbb{I}_{2}\big(t,3/4,\lambda_{N,1}^{2}/2,|K_{4}|^{5}+\left\|\mathscr{O}^{\tau,N}\right\|_{l^{6}_{N}}^{9}\big)\right),$
@@
$K_{4}(t):=\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left(1+\left\|\mathscr{O}^{\tau,N}_{s}\right\|^{8}_{l_{N}^{\infty}}+\left\|\left(-A_{N}\right)^{\frac{1}{2}}\mathscr{O}^{\tau,N}_{s}\right\|^{4}_{l_{N}^{4}}+\mathscr{R}_{s}^{8}+\tau^{-\frac{1}{3}}\left\|\mathscr{O}_{s}^{\tau,N}-\mathscr{O}_{\lfloor s\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{2}+\left\|\mathscr{O}_{s}^{\tau,N}-\mathscr{O}_{\lfloor s\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{4}\right)\mathrm{d}s.$
@@
$\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|u^{\tau,N}(t_{m},\cdot)\right\|_{l^{\infty}_{N}}^{p}\right]$ $\leq C\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|U_{t_{i}}^{\tau,N}\right\|_{w^{1,2}_{N}}^{p}\right]$
@@
which completes the proof.
4.2 Uniform strong convergence of the numerical scheme
To establish the uniform strong convergence of the numerical scheme, we require an assumption stronger than Assumption 2. {assumption} Suppose that the initial condition $u_{0}\in\mathcal{C}^{5}(\mathcal{O}).$ The main result of this subsection is given in the following theorem.
@@ 4.2 Uniform strong convergence of the numerical scheme
Suppose that Assumptions 2, 2, 4.2 hold and $h^{-1}\tau^{9}\leq 1$. Then for $p\geq 1$, there exists a constant $C>0$, independent of $\tau$ and $h$, such that
$\sup_{t\geq 0}\left\|u(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (74)
@@
$u^{N}(t,x)=P_{N}U^{N}_{t}(x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta_{N}G_{t-s}^{N}(x,y)f(u^{N}\left(s,\kappa_{N}(y)\right))\mathrm{d}y\mathrm{d}s$ (76)
@@
$\left\|u(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}\leq$ $\left\|u(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}+\left\|u^{N}(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}.$ (77)
Following the methods used to prove Theorems 3.1 and 4.2 in Deng& Cao (2025), we can replicate those steps with only minor modifications to establish the next two propositions. Then Theorem 4.7 is a straightforward conclusion of two auxiliary results stated below.
@@
Suppose that Assumptions 2, 2, 4.2 hold. Then for $p\geq 1$, there exists a constant $C>0$, independent of $N$, such that
$\sup_{t\geq 0}\left\|u(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon},$ (78)
@@
Suppose that Assumptions 2, 2, 4.2 hold and $h^{-1}\tau^{9}\leq 1$. Then for $p\geq 1$, there exists a constant $C>0$, independent of $\tau$ and $h$, such that
$\sup_{t\geq 0}\left\|u^{N}(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (79)
@@
The proof of this proposition is similar to Theorem 5.7 in Deng& Cao (2025). Therefore, these detailed steps are omitted.
5 Approximation of invariant measures
5.1 Existence and uniqueness of invariant measure for the numerical solution
Since $u^{\tau,N}(t_{m},\cdot)$ is defined as the interpolation of $U^{\tau,N}_{t_{m}}$, the existence and uniqueness of the invariant measure for $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ on $V^{N}$ is equivalent to that for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ on $\mathbb{R}^{N}$, where $V^{N}=\text{Span}\left\{P_{N}\phi_{N,0},P_{N}\phi_{N,1},\cdots,P_{N}\phi_{N,N-1}\right\}$. We develop an new ergodicity framework, building on the work of Mattingly, Stuart & Higham (2002), to prove the existence and uniqueness of the invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$. This guarantees the corresponding result for $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ . We are aware of several notable works, such as Chen et al. (2017, 2020), which applied this theory to establish the existence and uniqueness of invariant measures for numerical approximations to SPDEs. These studies primarily focused on the numerical invariant measures for SPDEs under Dirichlet boundary conditions. We aim to generalize the application of this theory to the study of numerical invariant measures for SPDEs subject to Neumann boundary conditions. The key point is to establish a minorization condition restricted to the hyperplane $D_{\alpha}:=\left\{\vec{u}\in\mathbb{R}^{N}:\ \ \frac{1}{\pi}\left\langle\vec{u},\phi_{N,0}\right\rangle_{l^{2}_{N}}=\alpha\right\}$. Hence we introduce the minorization condition restricted to $D_{\alpha}$.
{assumption}

### v3

@@ 4.  Fully discrete scheme and uniform strong convergence
$\left\{\begin{aligned} &\mathrm{d}V^{\tau,N}_{t}+A_{N}^{2}V^{\tau,N}_{t}\mathrm{d}t=A_{N}e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor})\mathrm{d}t,\quad t>0,\\ &V^{\tau,N}_{0}=u_{0}^{N}.\end{aligned}\right$. (4.17)
Taking the inner product on (4.17) with $V^{\tau,N}_{t}$, we obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\big\|V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}\right)=$ $-\big\|A_{N}V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}+\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor}),A_{N}\big(e^{-A_{N}^{2}(t-\lfloor t\rfloor)}-I\big)V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (4.18)
@@
Using Hölder’s inequality, the estimate (4.7) with $\gamma_{1}=\frac{1}{2}$, $\gamma_{2}=1$, and $\gamma_{3}=\frac{1}{4}$, (4.4) and Young’s inequality, we obtain
$\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor}),A_{N}\big(I-e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\big)V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (4.19)
@@
In order to estimate the third term on the right-hand side of (4.18), we need to establish the continuity of $U^{\tau,N}_{t}$. We rewrite (4.8) as
@@
Combining this with (4.6), we obtain
$\big\|U_{t}^{\tau,N}\big\|_{w^{1,2}_{N}}\leq$ $\left\|e^{-A_{N}^{2}(t-\lfloor t\rfloor)}U_{\lfloor t\rfloor}^{\tau,N}\right\|_{w^{1,2}_{N}}+(t-\lfloor t\rfloor)\left\|A_{N}e^{-A_{N}^{2}(t-\lfloor t\rfloor)}\tilde{F}_{N}\big(U_{\lfloor t\rfloor}^{\tau,N}\big)\right\|_{w^{1,2}_{N}}$ (4.20)
@@
where $\mathscr{R}_{t}:=\big\|\mathscr{O}_{t}^{\tau,N}\big\|_{w^{1,2}_{N}}+\big\|\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}$. Utilizing (4.8), we can deduce
$\left\|U_{t}^{\tau,N}-U_{\lfloor t\rfloor}^{\tau,N}\right\|_{l_{N}^{2}}$ (4.21)
@@
By using (4.7) with $\gamma_{1}=0,\ \gamma_{2}=1$ and $\gamma_{3}=\frac{1}{2}$, and Assumption 2.3, we have
$\mathscr{K}_{1}=\left\|e^{-A_{N}^{2}\lfloor t\rfloor}\left(e^{-A_{N}^{2}(t-\lfloor t\rfloor)}-I\right)u_{0}^{N}\right\|_{l^{2}_{N}}\leq C(t-\lfloor t\rfloor)^{\frac{1}{2}}\left\|A_{N}u_{0}^{N}\right\|_{l^{2}_{N}}\leq C\tau^{\frac{1}{2}}.$ (4.22)
It follows from (4.7) with $\gamma_{1}=1,\ \gamma_{2}=\frac{1}{2}$ and $\gamma_{3}=\frac{1}{2}$ that
$\mathscr{K}_{2}\leq$ $C\int_{0}^{\lfloor t\rfloor}(\lfloor t\rfloor-\lfloor s\rfloor)^{-\frac{3}{4}}(t-\lfloor t\rfloor)^{\frac{1}{2}}e^{-\frac{1}{2}(\lfloor t\rfloor-\lfloor s\rfloor)}\left\|(-A_{N})^{\frac{1}{2}}\tilde{F}_{N}\big(U_{\lfloor s\rfloor}^{\tau,N}\big)\right\|_{l^{2}_{N}}\mathrm{d}s$ (4.23)
@@
Similarly, by (4.7) and (4.4) we can obtain
$\mathscr{K}_{3}\leq\frac{(t-\lfloor t\rfloor)^{\frac{3}{4}}\left(1+\big\|U_{\lfloor s\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}^{3}\right)}{1+\tau\big\|U_{\lfloor s\rfloor}^{\tau,N}\big\|_{w^{1,2}_{N}}^{12}}\leq C\tau^{\frac{1}{2}}.$ (4.24)
By substituting (4.22)-(4.24) into (4.21), we obtain
$\left\|U_{s}^{M,N}-U_{\lfloor s\rfloor}^{M,N}\right\|_{l^{2}_{N}}\leq$ $C\tau^{\frac{1}{4}}+\left\|\mathscr{O}_{s}^{\tau,N}-\mathscr{O}_{\lfloor s\rfloor}^{\tau,N}\right\|_{l^{2}_{N}},$ (4.25)
which combined with Young’s inequality, (2.6), (4.3) and (4.20) yields
$\big\langle\tilde{F}_{N}(U^{\tau,N}_{\lfloor t\rfloor})-\tilde{F}_{N}(U^{\tau,N}_{t}),A_{N}V^{\tau,N}_{t}\big\rangle_{l^{2}_{N}}$ (4.26)
@@
Then it follows from (4.18), (4.19) and (4.26)
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left(\big\|V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}\right)=$ $-\frac{3}{4}\big\|A_{N}V^{\tau,N}_{t}\big\|^{2}_{l^{2}_{N}}+C\left(1+\mathscr{R}_{t}^{8}+\tau^{-\frac{1}{3}}\left\|\mathscr{O}_{t}^{\tau,N}-\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{2}+\left\|\mathscr{O}_{t}^{\tau,N}-\mathscr{O}_{\lfloor t\rfloor}^{\tau,N}\right\|_{l^{2}_{N}}^{4}\right)$ (4.27)
@@
Similar to the treatment of (3.12), we can obtain
@@
$\leq$ $C\Big(1+\|u^{N}_{0}\|_{l^{\infty}_{N}}^{20}+\left\|(-A_{N})^{\frac{1}{2}}u^{N}_{0}\right\|_{l^{2}_{N}}^{10}+\mathbb{I}_{1}\big(t,3/4,\lambda_{N,1}^{2}/2,|K_{4}|^{3}+\left\|\mathscr{O}^{\tau,N}\right\|_{l^{6}_{N}}^{3}\big)$
$\qquad+\mathbb{I}_{2}\big(t,3/4,\lambda_{N,1}^{2}/2,|K_{4}|^{5}+\left\|\mathscr{O}^{\tau,N}\right\|_{l^{6}_{N}}^{9}\big)\Big),$
@@
$K_{4}(t)$= $\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\Big(1+\|\mathscr{O}^{\tau,N}_{s}\|_{l_{N}^{\infty}}^{8}+\|(-A_{N})^{\frac{1}{2}}\mathscr{O}^{\tau,N}_{s}\|_{l_{N}^{4}}^{4}+\mathscr{R}_{s}^{8}$
$\qquad+\tau^{-1/3}\|\mathscr{O}^{\tau,N}_{s}-\mathscr{O}^{\tau,N}_{\lfloor s\rfloor}\|_{l_{N}^{2}}^{2}+\|\mathscr{O}^{\tau,N}_{s}-\mathscr{O}^{\tau,N}_{\lfloor s\rfloor}\|_{l_{N}^{2}}^{4}\Big)\,ds.$
@@
$\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|u^{\tau,N}(t_{m},\cdot)\right\|_{l^{\infty}_{N}}^{p}\right]$
$\leq$ $C\sup_{\tau\in(0,1),N\in\mathbb{N}}\sup_{i\geq 0}\mathbb{E}\left[\left\|U_{t_{i}}^{\tau,N}\right\|_{w^{1,2}_{N}}^{p}\right]$
@@
which completes the proof. ∎
4.Uniform strong convergence of the numerical scheme
To establish the uniform strong convergence of the numerical scheme, we require an assumption stronger than Assumption 2.3.
Assumption 4.5.
Suppose that the initial condition $u_{0}\in\mathcal{C}^{5}(\mathcal{O}).$
The main result of this subsection is given in the following theorem.
@@
Suppose that Assumptions 2.2, 2.4, 4.5 hold and $h^{-1}\tau^{9}\leq 1$. Then for $p\geq 1$, there exists a constant $C>0$, independent of $\tau$ and $h$, such that
$\sup_{t\geq 0}\left\|u(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (4.28)
@@
$u^{N}(t,x)=P_{N}U^{N}_{t}(x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta_{N}G_{t-s}^{N}(x,y)f(u^{N}\left(s,\kappa_{N}(y)\right))\mathrm{d}y\mathrm{d}s$ (4.30)
@@
$\left\|u(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}\leq$ $\left\|u(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}+\left\|u^{N}(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}\left(\Omega;L^{\infty}\right)}.$ (4.31)
Following the methods used to prove Theorems 3.1 and 4.2 in [19], we can replicate those steps with only minor modifications to establish the next two propositions. Then Theorem 4.2 is a straightforward conclusion of two auxiliary results stated below.
@@
Suppose that Assumptions 2.2, 2.4, 4.5 hold. Then for $p\geq 1$, there exists a constant $C>0$, independent of $N$, such that
$\sup_{t\geq 0}\left\|u(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon},$ (4.32)
@@
Suppose that Assumptions 2.2, 2.4, 4.5 hold and $h^{-1}\tau^{9}\leq 1$. Then for $p\geq 1$, there exists a constant $C>0$, independent of $\tau$ and $h$, such that
$\sup_{t\geq 0}\left\|u^{N}(t,\cdot)-u^{\tau,N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq C\left(h^{1-\epsilon}+\tau^{\frac{3}{8}-\frac{\epsilon}{2}}\right),$ (4.33)
@@
The proof of this proposition is similar to Theorem 5.7 in [19]. Therefore, these detailed steps are omitted.
5.  Approximation of invariant measures
5.Existence and uniqueness of invariant measure for the numerical solution
Since $u^{\tau,N}(t_{m},\cdot)$ is defined as the interpolation of $U^{\tau,N}_{t_{m}}$, the existence and uniqueness of the invariant measure for $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ on $V^{N}$ is equivalent to that for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$ on $\mathbb{R}^{N}$, where $V^{N}=\text{Span}\left\{P_{N}\phi_{N,0},P_{N}\phi_{N,1},\cdots,P_{N}\phi_{N,N-1}\right\}$. We develop an new ergodicity framework, building on the work of [31], to prove the existence and uniqueness of the invariant measure for $\left\{U^{\tau,N}_{t_{m}}\right\}_{m\in\mathbb{N}}$. This guarantees the corresponding result for $\left\{u^{\tau,N}(t_{m},\cdot)\right\}_{m\in\mathbb{N}}$ . We are aware of several notable works, such as [9, 10], which applied this theory to establish the existence and uniqueness of invariant measures for numerical approximations to SPDEs. These studies primarily focused on the numerical invariant measures for SPDEs under Dirichlet boundary conditions. We aim to generalize the application of this theory to the study of numerical invariant measures for SPDEs subject to Neumann boundary conditions. The key point is to establish a minorization condition restricted to the hyperplane $D_{\alpha}:=\left\{\vec{u}\in\mathbb{R}^{N}:\ \ \frac{1}{\pi}\left\langle\vec{u},\phi_{N,0}\right\rangle_{l^{2}_{N}}=\alpha\right\}$. Hence we introduce the minorization condition restricted to $D_{\alpha}$.
Assumption 5.1.